OpenStudy (anonymous):

write an equation in slope intercept form for each line according to the given information. Contains (0,6) Is parallel to 4y = 20x - 3 Show all work.

hartnn (hartnn):

can you find slope of 4y = 20x - 3 ?

OpenStudy (anonymous):

y=5x-3/4

OpenStudy (anonymous):

i think

hartnn (hartnn):

yes, so compare that with y=mx+c them slope =m=...?

OpenStudy (anonymous):

5?

hartnn (hartnn):

thats correct, now parallel lines have same slope, so, your required line also has slope of 5 also, since you have point 0,6 on it, can you find the equation of line in slope point form, \(y-y_1=m(x-x_1)\)

hartnn (hartnn):

m=5, x1 =0,y1 = 6 just plug in values!

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^